OpenStudy (anonymous):

Solve V = π r2h for h. Show all work. Help?

OpenStudy (saifoo.khan):

BONEZ.

OpenStudy (anonymous):

Saifoo

OpenStudy (anonymous):

\[h = \frac{V}{2 \pi r}\]

OpenStudy (chaise):

V = π r2h H = v/(pi*r*2) Need help on how to do this?

OpenStudy (saifoo.khan):